Find the total number of nucleons present in 12g of 12C atoms

Grade:12

2 Answers

Sakshi
8 years ago
In 12 g of C-12, there are (6+6 protons + neutrons) * 6.022*1023
= 12*6.022*1023 nucleons
Anjali kumari
5 years ago
As mole=given mass/molecular mass =12/12=1 mole.Now total number of nucleons of 1 mole carbon=6 mole proton+6 mole neutron=12 moles.